# Current address: ]Susi-Nicoletti-Weg 9, Vienna A-1100, Austria

### v1

@@ IV.1 Modelling mixed states with pure states
where we defined $|\psi^{\prime}_{S}\rangle=\mathrm{BS}\cdot|\psi_{S}\rangle$, $|\psi^{\prime}_{I}\rangle=\mathrm{DM}_{I}\cdot|\psi_{I}\rangle$ and $|\Psi^{\prime}\rangle=\mathrm{BS}\cdot\mathrm{DM}_{I}\cdot|\Psi\rangle$. Remember that $\langle\psi_{w}|\psi^{\prime}_{I}\rangle=\langle\psi_{w}|\Psi\rangle=0$.
@@
Before we proceed to discuss the visibilities of these detection probabilities, let us emphasize that there is not a single term in Eq. (37) that would contain inner products of $|\psi^{\prime}_{I}\rangle$ with itself or its projections onto some subspace. For that reason, there is no term that would contain the inner product $\langle e_{H}|e_{V}\rangle$ (32), whose determination is necessary for the characterization of the mixed state $\hat{\rho}_{I}$. This observation is independent of particular forms of $|\psi^{\prime}_{S}\rangle$ and $|\Psi^{\prime}\rangle$ and is thus quite general. We, therefore, conclude with the important statement that the imbalanced ZWM interferometer in Fig. 2 does not allow for the perfect reconstruction of mixed states unless additional constraints are enforced. We come back to this problem in Sec. IV.4.
IV.2 Visibilities
In this section, we carry out the derivation of visibilities in different bases in a way analogous to that in Sec. III. When the signal photon’s polarization is measured in an orthonormal basis $\{|k\rangle,|k^{\perp}\rangle\}$, the probability of detecting $|k\rangle$ is given in Eq. (37), which profile exhibits interference as the phase $\phi$ is varied. The visibility of the emergent interference pattern is given by (see Appendix B)
@@ IV.2 Visibilities
where $c_{k}$ and $z_{k}$ are given in Eqs. (38) and (39), respectively. This expression attains a conveniently simple form when we set $P=1$ and choose the signal photon’s state $|\psi^{\prime}_{S}\rangle$ to be unbiased with both $|k\rangle$ and $|k^{\perp}\rangle$, i.e., $|\langle\psi^{\prime}_{S}|k\rangle|=|\langle\psi^{\prime}_{S}|k^{\perp}\rangle|=1/\sqrt{2}$. It reduces to (see Appendix C)
@@
where $|k^{\ast}\rangle$ is the “complex conjugate” of vector $|k\rangle$ in the sense that if $|k\rangle=V|H\rangle$ for some unitary $V$, then $|k^{\ast}\rangle=V^{\ast}|H\rangle$, where the star stands for the complex conjugation. The visibilities are thus equal to the overlap between the idler state (with its environment) and the complex conjugate basis state with environment $|e_{\Psi}\rangle$.
The squares of visibilities show interesting properties. At first, note that from Eq. (41) and the definition of $|\psi^{\prime}_{I}\rangle$ it follows that
$=$ $\langle\psi_{I}|\,\hat{\mathcal{V}}_{k}\,|\psi_{I}\rangle,$
where we defined an operator
$\hat{\mathcal{V}}_{k}=T^{2}(\mathrm{DM}_{I}\cdot|k^{\ast}\rangle\langle k^{\ast}|\cdot\mathrm{DM}_{I}^{\dagger})\otimes|e_{\Psi}\rangle\langle e_{\Psi}|.$ (43)
This visibility operator deserves more attention and is further studied in Sec. V. For the other basis vector $|k^{\perp}\rangle$ we analogously obtain
$\mathcal{V}_{k^{\perp}}^{2}=\langle\psi_{I}|\,(T^{2}(\mathbb{1}\otimes|e_{\Psi}\rangle\langle e_{\Psi}|)-\hat{\mathcal{V}}_{k})\,|\psi_{I}\rangle,$ (44)
which evaluates to
$\mathcal{V}_{k^{\perp}}^{2}=T^{2}(\alpha^{2}m_{H}^{2}+\beta^{2}m_{V}^{2})-\mathcal{V}_{k}^{2},$ (45)
where we defined
$m_{H}=|\langle e_{H}|e_{\Psi}\rangle|,\quad m_{V}=|\langle e_{V}|e_{\Psi}\rangle|.$ (46)
These parameters quantify the coherence of the $H$ and $V$ modes of the idler state with the second source Q2 and their properties are further studied in Sec. IV.4. From the above formulas, it directly follows that the sum of visibilities for the two orthogonal vectors is constant. This constant quantifies the level of coherence between the first and the second source, and we refer to it as the zeroth visibility Stokes parameter $\mathcal{S}_{0}$, given by
$\mathcal{S}_{0}\equiv\mathcal{V}_{k}^{2}+\mathcal{V}_{k^{\perp}}^{2}=T^{2}(\alpha^{2}m_{H}^{2}+\beta^{2}m_{V}^{2}).$ (47)
It is easy to see that $0\leq\mathcal{S}_{0}\leq T^{2}\leq 1$. For the difference of the two visibilities, we then obtain
$\mathcal{V}_{k}^{2}-\mathcal{V}_{k^{\perp}}^{2}=2\mathcal{V}_{k}^{2}-\mathcal{S}_{0}.$ (48)
This expression plays a crucial role and is discussed in the next section.
For completeness, let us list the visibilities of detection probabilities for the typical choice of single-qubit bases H/V, D/A, and R/L. For these, we get $(k,k^{\ast})\in\{(H,H),(V,V),(D,D),(A,A),(L,R),(R,L)\}$ and the visibilities explicitly read
@@
Compare these expressions with Eqs. (19)–(21).
IV.3 Visibility Stokes parameters
We can define the visibility Stokes parameters for mixed states in the exact same way as we did for pure states in Eqs. (28)–(30), where the visibilities are now given by Eq. (41). The visibility Stokes parameters then explicitly read
@@ IV.3 Visibility Stokes parameters
$\mathcal{S}_{z}$ $=$ $T^{2}\,(\alpha^{2}\,m^{2}_{H}-\beta^{2}\,m^{2}_{V}),$ (54)
where we defined $\Delta\varphi=\arg\left(\langle e_{H}|e_{\Psi}\rangle\right)-\arg\left(\langle e_{V}|e_{\Psi}\rangle\right)$. The transmission coefficient $T$ can be measured for a given experimental setup independently of a particular state of the idler photon, and one can thus adjust the formulas above by removing $T$. For this reason, from now on, we effectively set $T=1$.
@@
$\vec{\mathcal{S}}=(\mathcal{S}_{x},\mathcal{S}_{y},\mathcal{S}_{z}).$ (55)
$\mathcal{S}_{x}^{2}+\mathcal{S}_{y}^{2}+\mathcal{S}_{z}^{2}=\mathcal{S}_{0}^{2}$ (56)
and so the norm of the visibility Bloch vector is equal to the zeroth parameter: $|\vec{\mathcal{S}}|=\mathcal{S}_{0}$. An important question is how close this visibility Bloch vector is to the actual Bloch vector $\vec{r}=(x,y,z)$. It is straightforward to show from the explicit expressions in Eqs. (12)–(14) and Eqs. (52)–(54) together with the inequality in Eq. (59), discussed in the next section, that
$\mathscr{P}\leq(\vec{r}\cdot\vec{\mathcal{S}})+1-\mathcal{S}_{0}.$ (58)
When the inequality (58) is understood as a constraint on varying visibility Stokes parameters while the Bloch vector $\vec{r}$ is fixed, it turns out that this inequality represents a rotational ellipsoid, see Appendix D. For a given polarization state of a photon, one can thus, in principle, obtain many visibility Bloch vectors, depending on coherence conditions, and all these vectors form an ellipsoid, see Fig. 4.
Several other algebraic properties of the visibility Stokes parameters can be found in Appendix D. In the next section, the aforementioned issues are discussed in detail, and possible solutions to the ambiguity of the visibility Stokes parameters are proposed.
IV.4 Role of asymmetric coherence
The quantities $q$, $m_{H}$, $m_{V}$, and $\Delta\varphi$ defined in the previous section are not completely independent of each other. As shown in Appendix E, they have to satisfy inequality
@@ IV.4 Role of asymmetric coherence
Another notable situation is when only the vertical polarization interacts with the environment. In contrast, the horizontal polarization is not affected, as is the case in some birefringent materials, and so $|e_{H}\rangle=|e_{\Psi}\rangle$ with $m_{H}=1$. For such values, the inequality (59) can be satisfied only when $m_{V}=q$ and $\Delta\varphi=0$. This situation is depicted as a red line in Fig. 5. Assuming that $T=1$, the visibility Stokes parameters in Eqs. (47) and (52)–(54) reduce to
@@
The first two formulas are identical to the real Bloch vector coordinates in Eqs. (12) and (13). The third formula does not match the $z$-coordinate in Eq. (14), but we can remedy this easily by noting that due to the normalization $\alpha^{2}+\beta^{2}=1$, the $z$-coordinate is equal to
@@
In this scenario, we can thus reconstruct the polarization state of the idler photon. This special case was considered in Ref. [6], where the constraint $m_{H}=1$ was enforced by the calibration of the setup performed before the tomography itself. Obviously, when $m_{V}=1$, then analogously $q=m_{H}$ and a similar discussion can be made.
@@
from where we calculate $q$. At this point, we can renormalize expressions (66), (67), and (68) by dividing by $\mathcal{S}_{0}$ and multiplying by $q$ to obtain the real Bloch coordinates of the mixed state $\hat{\rho}_{I}$ in Eq.  (11). This scenario is depicted as a blue line in Fig. 5.
V Measurement operators corresponding to visibility measurements
The standard Stokes parameters can be expressed as expectation values of measurement operators in a given quantum state. In this section, we analogously introduce the operators corresponding to the visibility Stokes parameters. As follows from Eq. (42), one can express the visibilities as expectation values of specific measurement operators defined in Eq. (43). These can be used to introduce operators that correspond to the visibility Stokes parameters
@@ V Measurement operators corresponding to visibility measurements
Even though the operators $\hat{\mathcal{S}}_{x}$, $\hat{\mathcal{S}}_{y}$, and $\hat{\mathcal{S}}_{z}$ derived above resemble the Pauli operators in Eqs. (6)–(8), it is important to emphasize that they do not form a complete measurement in the usual sense. Formally speaking, each visibility Stokes operator $\hat{\mathcal{S}}_{k}$ corresponds to two positive measurement operators $\hat{\mathcal{V}}_{k}$ and $\hat{\mathcal{V}}_{k^{\perp}}$. These two operators nevertheless do not satisfy the completeness relation $\sum_{k}\hat{\mathcal{V}}_{k}=\mathbb{1}$. Instead, they sum up to $\mathcal{S}_{0}$, cf. Eq. (47). This can be remedied by expanding each set of visibility operators with an operator $\hat{\mathcal{V}}_{\mathrm{incoh}}\equiv\mathbb{1}-\hat{\mathcal{S}}_{0}$, quantifying the incoherence between the two sources. Unlike the standard polarization measurement, where e.g. $\hat{\Pi}_{H}+\hat{\Pi}_{V}=\mathbb{1}$ with $\hat{\Pi}_{k}=|k\rangle\langle k|$, here we have $\hat{\mathcal{V}}_{H}+\hat{\mathcal{V}}_{V}+\hat{\mathcal{V}}_{\mathrm{incoh}}=\mathbb{1}$. Let us note that for $T=1$, the visibility operators, including $\hat{\mathcal{V}}_{\mathrm{incoh}}$ are projectors.
VI Post-measurement states
@@
The final state is a uniform mixture of the maximally mixed state $\mathbb{1}/2$ and the original pure state $|\psi^{\prime}_{I}\rangle$ in path $c$ (15). Evidently, the post-measurement state of the idler photon still contains some information about its original state of polarization. Also, note that the post-measurement state is not an eigenstate of a Pauli matrix as one would expect in the standard polarization tomography.
VII Conclusion
@@ Appendix C Visibilities for unbiased signal states
$\langle k|\Phi^{+}\rangle$ $=\langle H|(V^{\dagger}\otimes\mathbb{1})|\Phi^{+}\rangle$
$=\langle H|(\mathbb{1}\otimes V^{\ast})(V^{\dagger}\otimes V^{T})|\Phi^{+}\rangle$
$=\langle H|(\mathbb{1}\otimes V^{\ast})|\Phi^{+}\rangle$
$=(1/\sqrt{2})V^{\ast}|H\rangle,$ (86)
where $|k\rangle=V|H\rangle$ for a specific unitary $V$, and where we used the unitary invariance for $U=V^{\dagger}$.
We model the symmetric beam splitter by a Hadamard matrix. If we fix the projector $\hat{\Pi}_{j}$ to act on path $b$, it is easy to show that the brakets in the definition of $c_{j}$ in Eq. (38) reduce to $\langle\psi^{\prime}_{S}|\hat{\Pi}_{j}|\psi^{\prime}_{S}\rangle=\langle\psi_{S}|\mathrm{BS}^{\dagger}\hat{\Pi}_{j}\mathrm{BS}|\psi_{S}\rangle=(1/2)\langle\psi_{S}|\hat{\Pi}_{j}|\psi_{S}\rangle$ and $\langle\Psi^{\prime}|\hat{\Pi}_{j}|\Psi^{\prime}\rangle=\langle\Phi^{+}|\mathrm{BS}^{\dagger}\hat{\Pi}_{j}\mathrm{BS}|\Phi^{+}\rangle=(1/2)\langle\Phi^{+}|\hat{\Pi}_{j}|\Phi^{+}\rangle$. Similarly, $z_{j}$ in Eq. (39) reduces to $\langle\psi^{\prime}_{S}|\otimes\langle\psi^{\prime}_{I}|\hat{\Pi}_{j}|\Psi^{\prime}\rangle=\langle\psi_{S}|\mathrm{BS}^{\dagger}\hat{\Pi}_{j}\mathrm{BS}\otimes\langle\psi_{I}|\Psi\rangle=(-1/2)\langle\psi_{S}|\hat{\Pi}_{j}\otimes\langle\psi_{I}|\Psi\rangle$.
From Eq. (86) we further get $\langle\Phi^{+}|\hat{\Pi}_{j}|\Phi^{+}\rangle=\langle\Phi^{+}|j\rangle\langle j|\Phi^{+}\rangle=1/2$. As mentioned in the main text, we set the signal’s state $|\psi_{S}\rangle$ to be mutually unbiased with the projection vector $|j\rangle$, where $\hat{\Pi}_{j}=|j\rangle\langle j|$, that is $\langle\psi_{S}|j\rangle=(1/\sqrt{2})\exp(i\varphi_{j})$ for some real $\varphi_{j}$. From there it follows that $\langle\psi_{S}|\hat{\Pi}_{j}|\psi_{S}\rangle=1/2$ and $\langle\psi_{S}|\hat{\Pi}_{j}\otimes\langle\psi_{I}|\Psi\rangle=(1/\sqrt{2})\exp(i\varphi_{j})\langle j|\langle\psi_{I}|\Phi^{+}\rangle|e_{\Psi}\rangle=(1/2)\exp(i\varphi_{j})\langle\psi_{I}|V_{j}^{\ast}|H\rangle|e_{\Psi}\rangle$, where $|j\rangle=V_{j}|H\rangle$ and where we used Eq. (86).
Taken all these simplifications into account, we can recast the form of $c_{j}$ and $|z_{j}|$ into $c_{j}=(1/4)(1+P^{2})$ and $|z_{j}|=(1/4)|\langle\psi_{I}|j^{\ast}\rangle|e_{\Psi}\rangle|$, where we define $|j^{\ast}\rangle=V_{j}^{\ast}|H\rangle$. The ratio of these two quantities determines the visibility, for which we obtain
@@
Appendix D Relations of visibility Stokes parameters

### v2

@@ IV.2 Photon evolution through setup
where we defined $|\psi^{\prime}_{S}\rangle=\mathrm{BS}\cdot|\psi_{S}\rangle$, $|\psi^{\prime}_{I}\rangle=\mathrm{DM}_{I}\cdot|\psi_{I}\rangle$ and $|\Psi^{\prime}\rangle=\mathrm{BS}\cdot\mathrm{DM}_{I}\cdot|\Psi\rangle$. Remember that $\langle\psi_{w}|\psi^{\prime}_{I}\rangle=\langle\psi_{w}|\Psi^{\prime}\rangle=0$.
@@
Before we proceed to discuss the visibilities of these detection probabilities, let us emphasize that there is not a single term in Eq. (38) that would contain inner products of $|\psi^{\prime}_{I}\rangle$ with itself or its projections onto some subspace. For that reason, there is no term that would contain the inner product $\langle e_{H}|e_{V}\rangle$ (33), whose determination is necessary for the characterization of the mixed state $\hat{\rho}_{I}$. This observation is independent of particular forms of $|\psi^{\prime}_{S}\rangle$ and $|\Psi^{\prime}\rangle$ and is thus quite general. We, therefore, conclude with the important statement that the imbalanced ZWM interferometer in Fig. 2 does not allow for the perfect reconstruction of mixed states unless additional constraints are enforced. We come back to this problem in Sec. IV.5.
IV.3 Visibilities
In this section, we carry out the derivation of visibilities in different bases in a way analogous to that in Sec. III.2. When the signal photon’s polarization is measured in an orthonormal basis $\{|k\rangle,|k^{\perp}\rangle\}$, the probability of detecting $|k\rangle$ is given in Eq. (38), which profile exhibits interference as the phase $\phi$ is varied. The visibility of the emergent interference pattern is given by (see Appendix B)
@@ IV.3 Visibilities
where $c_{k}$ and $z_{k}$ are defined in Eqs. (39) and (40), respectively. This expression attains a conveniently simple form when we set $P=1$ and choose the signal photon’s state $|\psi_{S}\rangle$ to be unbiased with both $|k\rangle$ and $|k^{\perp}\rangle$, i.e., $|\langle\psi_{S}|k\rangle|=|\langle\psi_{S}|k^{\perp}\rangle|=1/\sqrt{2}$. It reduces to (see Appendix C)
@@
where $|k^{\ast}\rangle$ is the “complex conjugate” of vector $|k\rangle$ in the sense that if $|k\rangle=U|H\rangle$ for some unitary $U$, then $|k^{\ast}\rangle=U^{\ast}|H\rangle$, where the star stands for the complex conjugation. The visibilities are thus equal to the overlap between the idler state (with its environment) and the complex conjugate basis state (with environment $|e_{\Psi}\rangle$).
For completeness, let us list the visibilities of detection probabilities for the typical choice of single-qubit bases $H/V$, $D/A$, and $L/R$. For these, we get $(k,k^{\ast})\in\{(H,H),(V,V),(D,D),(A,A),(L,R),(R,L)\}$ and the visibilities explicitly read
@@
Compare these expressions with Eqs. (20)–(22). The products of environmental states apparently play an important role and for convenience we introduce the following notation
$m_{H}$ $=|\langle e_{H}|e_{\Psi}\rangle|,\quad m_{V}=|\langle e_{V}|e_{\Psi}\rangle|,$ (46)
$\Delta\varphi$ $=\arg\left(\langle e_{H}|e_{\Psi}\rangle\right)-\arg\left(\langle e_{V}|e_{\Psi}\rangle\right).$ (47)
These parameters quantify the coherence of the $H$ and $V$ modes of the idler state with the second source Q2 and their properties are further studied in Sec. IV.5.
IV.4 Visibility Stokes parameters for mixed states
In the previous section, formulas (43)–(45) connect the form of environmental states and the visibilities of signal photon detection rates in individual Pauli bases. It is easy to verify from these expressions that
$\mathcal{V}_{H}^{2}+\mathcal{V}_{V}^{2}=\mathcal{V}_{D}^{2}+\mathcal{V}_{A}^{2}=\mathcal{V}_{L}^{2}+\mathcal{V}_{R}^{2}=\mathcal{S}_{0},$ (48)
where we defined
$\mathcal{S}_{0}\equiv T^{2}(\alpha^{2}m_{H}^{2}+\beta^{2}m_{V}^{2}).$ (49)
The sum of squares of visibilities in the two orthogonal states is thus constant. This constant quantifies the level of coherence between the first and the second source, and we refer to it as the zeroth visibility Stokes parameter $\mathcal{S}_{0}$. It is easy to check that $0\leq\mathcal{S}_{0}\leq T^{2}\leq 1$ (compare with Eq. (28)). In Appendix D we show that $\mathcal{V}_{k}^{2}+\mathcal{V}_{k^{\perp}}^{2}=\mathcal{S}_{0}$ for any orthonormal basis $\{|k\rangle,|k^{\perp}\rangle\}$, not just the Pauli bases.
We can define the other visibility Stokes parameters for mixed states in the exact same way as we did for pure states in Eqs. (29)–(31) except that now the visibilities are given by Eq. (42). The visibility Stokes parameters then explicitly read
@@ IV.4 Visibility Stokes parameters for mixed states
$\mathcal{S}_{z}$ $=$ $T^{2}\,(\alpha^{2}\,m^{2}_{H}-\beta^{2}\,m^{2}_{V}).$ (52)
The transmission coefficient $T$ can be measured for a given experimental setup independently of a particular state of the idler photon, and one can thus adjust the formulas above by removing $T$. For this reason, from now on, we effectively set $T=1$.
@@
In the next section, the aforementioned issues are discussed in detail, and possible solutions to the ambiguity of the visibility Stokes parameters are proposed.
IV.5 Role of asymmetric coherence
The quantities $q$, $m_{H}$, $m_{V}$, and $\Delta\varphi$ defined in the previous sections are not completely independent of each other. As shown in Appendix E, they have to satisfy the inequality
@@ IV.5 Role of asymmetric coherence
Another notable situation is when only the vertical polarization interacts with the environment. In contrast, the horizontal polarization is not affected, as is the case in some birefringent materials, and so $|e_{H}\rangle=|e_{\Psi}\rangle$ with $m_{H}=1$. For such values, the inequality (53) can be satisfied only when $m_{V}=q$ and $\Delta\varphi=0$. This situation is depicted as a red line in Fig. 3. Assuming that $T=1$, the visibility Stokes parameters in Eqs. (49)–(52) reduce to
@@
The first two formulas are identical to the standard Bloch vector coordinates in Eqs. (12) and (13). The third formula does not match the $z$-coordinate in Eq. (14), but we can remedy this easily by noting that due to the normalization $\alpha^{2}+\beta^{2}=1$, the $z$-coordinate is equal to
@@
In this scenario, we can thus reconstruct the polarization state of the idler photon. This special case was considered in Ref. [9], where the constraint $m_{H}=1$ was enforced. Obviously, when $m_{V}=1$, then analogously $q=m_{H}$ and a similar discussion can be made.
@@
from where we calculate $q$. At this point, we can renormalize expressions (60), (61), and (62) by dividing by $\mathcal{S}_{0}$ and multiplying by $q$ to obtain the standard Bloch coordinates of the mixed state $\hat{\rho}_{I}$ in Eq.  (11). This scenario is depicted as a blue line in Fig. 3.
V Algebraic properties of visibility Stokes parameters
$\vec{\mathcal{S}}=(\mathcal{S}_{x},\mathcal{S}_{y},\mathcal{S}_{z}).$ (64)
$\mathcal{S}_{x}^{2}+\mathcal{S}_{y}^{2}+\mathcal{S}_{z}^{2}=\mathcal{S}_{0}^{2}$ (65)
and so the norm of the visibility Bloch vector is equal to the zeroth parameter: $|\vec{\mathcal{S}}|=\mathcal{S}_{0}$. An important question is how close this visibility Bloch vector is to the actual Bloch vector $\vec{r}=(x,y,z)$. It is straightforward to show from the explicit expressions in Eqs. (12)–(14) and Eqs. (50)–(52) together with the inequality in Eq. (53), that
$\mathscr{P}\leq(\vec{r}\cdot\vec{\mathcal{S}})+1-\mathcal{S}_{0}.$ (67)
When the inequality (67) is understood as a constraint on varying visibility Stokes parameters while the Bloch vector $\vec{r}$ is fixed, it turns out that this inequality represents a rotational ellipsoid, see Appendix D. The ellipsoid has its center in point $\vec{r}/2$, its two loci are located in the origin $(0,0,0)$ and point $\vec{r}$, and its semiaxes read $1/2$ and $\sqrt{1-r^{2}}/2$. Depending on coherence conditions one can thus for a given polarization state of a photon obtain many visibility Bloch vectors and all these vectors form an ellipsoid, see Fig. 5.
Several other algebraic properties of the visibility Stokes parameters can be found in Appendix D.
VI Measurement operators corresponding to visibility measurements
The standard Stokes parameters can be expressed as expectation values of measurement operators in a given quantum state. In this section, we analogously introduce the operators corresponding to the visibility Stokes parameters. At first, note that from Eq. (42) and the definition of $|\psi^{\prime}_{I}\rangle$ it follows that
$=$ $\langle\psi_{I}|\,\hat{\mathcal{V}}_{k}\,|\psi_{I}\rangle,$
where we defined an operator
$\hat{\mathcal{V}}_{k}=T^{2}(\mathrm{DM}_{I}^{\dagger}\cdot|k^{\ast}\rangle\langle k^{\ast}|\cdot\mathrm{DM}_{I})\otimes|e_{\Psi}\rangle\langle e_{\Psi}|.$ (69)
One can therefore express the visibilities as expectation values of specific measurement operators. Following Eqs. (29)–(31), these operators can be used to introduce a triple of operators that correspond to the visibility Stokes parameters
@@ VI Measurement operators corresponding to visibility measurements
Even though the operators $\hat{\mathcal{S}}_{x}$, $\hat{\mathcal{S}}_{y}$, and $\hat{\mathcal{S}}_{z}$ derived above resemble the Pauli operators in Eqs. (6)–(8), it is important to emphasize that they do not form a complete measurement in the usual sense. Formally speaking, each visibility Stokes operator $\hat{\mathcal{S}}_{k}$ corresponds to two positive measurement operators $\hat{\mathcal{V}}_{k}$ and $\hat{\mathcal{V}}_{k^{\perp}}$. These two operators nevertheless do not satisfy the completeness relation $\sum_{k}\hat{\mathcal{V}}_{k}=\mathbb{1}$. Instead, they sum up to $\mathcal{S}_{0}$, cf. Eq. (48). This can be remedied by expanding each set of visibility operators with an operator $\hat{\mathcal{V}}_{\mathrm{incoh}}\equiv\mathbb{1}-\hat{\mathcal{S}}_{0}$, quantifying the incoherence between the two sources. Unlike the standard polarization measurement, where e.g. $\hat{\Pi}_{H}+\hat{\Pi}_{V}=\mathbb{1}$ with $\hat{\Pi}_{k}=|k\rangle\langle k|$, here we have $\hat{\mathcal{V}}_{H}+\hat{\mathcal{V}}_{V}+\hat{\mathcal{V}}_{\mathrm{incoh}}=\mathbb{1}$. Let us note that for $T=1$, the visibility operators, including $\hat{\mathcal{V}}_{\mathrm{incoh}}$, are projectors.
VII Post-measurement states
@@
The final state is a uniform mixture of the maximally mixed state $\mathbb{1}/2$ and the original pure state $|\psi^{\prime}_{I}\rangle$ in path $c$ (15). Evidently, the post-measurement state of the idler photon still contains some information about its original state of polarization.
VIII Conclusion
@@ Appendix C Visibilities for unbiased signal states
$\langle k|\Phi^{+}\rangle$ $=\langle H|(U^{\dagger}\otimes\mathbb{1})|\Phi^{+}\rangle$
$=\langle H|(\mathbb{1}\otimes U^{\ast})(U^{\dagger}\otimes U^{T})|\Phi^{+}\rangle$
$=\langle H|(\mathbb{1}\otimes U^{\ast})|\Phi^{+}\rangle$
$=(1/\sqrt{2})U^{\ast}|H\rangle,$ (86)
where $|k\rangle=U|H\rangle$ for a specific unitary $U$, and where we used the unitary invariance for $U^{\dagger}$.
We model the symmetric beam splitter by a Hadamard matrix. If we fix the projector $\hat{\Pi}_{j}$ to act (initially) on path $e$, it is easy to show that the brakets in the definition of $c_{j}$ in Eq. (39) reduce to $\langle\psi^{\prime}_{S}|\hat{\Pi}_{j}|\psi^{\prime}_{S}\rangle=\langle\psi_{S}|\mathrm{BS}^{\dagger}\hat{\Pi}_{j}\mathrm{BS}|\psi_{S}\rangle=(1/2)\langle\psi_{S}|\hat{\Pi}_{j}|\psi_{S}\rangle$ and $\langle\Psi^{\prime}|\hat{\Pi}_{j}|\Psi^{\prime}\rangle=\langle\Phi^{+}|\mathrm{BS}^{\dagger}\hat{\Pi}_{j}\mathrm{BS}|\Phi^{+}\rangle=(1/2)\langle\Phi^{+}|\hat{\Pi}_{j}|\Phi^{+}\rangle$. Similarly, $z_{j}$ in Eq. (40) reduces to $\langle\psi^{\prime}_{S}|\otimes\langle\psi^{\prime}_{I}|\hat{\Pi}_{j}|\Psi^{\prime}\rangle=\langle\psi_{S}|\mathrm{BS}^{\dagger}\hat{\Pi}_{j}\mathrm{BS}\otimes\langle\psi_{I}|\Psi\rangle=(-1/2)\langle\psi_{S}|\hat{\Pi}_{j}\otimes\langle\psi_{I}|\Psi\rangle$.
From Eq. (86) we further get $\langle\Phi^{+}|\hat{\Pi}_{j}|\Phi^{+}\rangle=\langle\Phi^{+}|j\rangle\langle j|\Phi^{+}\rangle=1/2$. As mentioned in the main text, we set the signal’s state $|\psi_{S}\rangle$ to be mutually unbiased with the projection vector $|j\rangle$, where $\hat{\Pi}_{j}=|j\rangle\langle j|$, that is $\langle\psi_{S}|j\rangle=(1/\sqrt{2})\exp(i\varphi_{j})$ for some real $\varphi_{j}$. From there it follows that $\langle\psi_{S}|\hat{\Pi}_{j}|\psi_{S}\rangle=1/2$ and $\langle\psi_{S}|\hat{\Pi}_{j}\otimes\langle\psi_{I}|\Psi\rangle=(1/\sqrt{2})\exp(i\varphi_{j})\langle j|\langle\psi_{I}|\Phi^{+}\rangle|e_{\Psi}\rangle=(1/2)\exp(i\varphi_{j})\langle\psi_{I}|U_{j}^{\ast}|H\rangle|e_{\Psi}\rangle$, where $|j\rangle=U_{j}|H\rangle$ and where we used Eq. (86).
Taken all these simplifications into account, we can recast the form of $c_{j}$ and $|z_{j}|$ into $c_{j}=(1/4)(1+P^{2})$ and $|z_{j}|=(1/4)|\langle\psi_{I}|j^{\ast}\rangle|e_{\Psi}\rangle|$, where we define $|j^{\ast}\rangle=U_{j}^{\ast}|H\rangle$. The ratio of these two quantities determines the visibility, for which we obtain
@@
Appendix D Notable properties of visibility Stokes parameters